# Monitoring AI-Modified Content at Scale: A Case Study on the Impact of ChatGPT on AI Conference Peer Reviews

### 11 Mar 2024 v1

@@ 2 Related Work
Text watermarking introduces a method to detect AI-generated text by embedding unique, algorithmically-detectable signals -known as watermarks- directly into the text. Early post-hoc approaches modify pre-existing text by leveraging synonym substitution (Chiang et al., 2003; Topkara et al., 2006b), syntactic structure restructuring (Atallah et al., 2001; Topkara et al., 2006a), or paraphrasing (Atallah et al., 2002). Increasingly, scholars have focused on integrating a watermark directly into an LLM’s decoding process. Kirchenbauer et al. (2023) split the vocabulary into red-green lists based on hash values of previous n-grams and then increase the logits of green tokens to embed the watermark. Zhao et al. (2023) use a global red-green list to enhance robustness. Hu et al. (2023b); Kuditipudi et al. (2023); Wu et al. (2023) study watermarks that preserve the original token probability distributions. Meanwhile, semantic watermarks (Hou et al., 2023; Fu et al., 2023; Liu et al., 2023) using input sequences to find semantically related tokens and multi-bit watermarks (Yoo et al., 2023; Fernandez et al., 2023) to embed more complex information have been proposed to improve certain conditional generation tasks. However, watermarking requires the involvement of the model or service owner, such as OpenAI, to implant the watermark. Concerns have also been raised regarding the potential for watermarking to degrade text generation quality and to compromise the coherence and depth of LLM responses (Singh & Zou, 2023). In contrast, our framework operates independently of the model or service owner’s intervention, allowing for the monitoring of AI-modified content without requiring their adoption.
@@ 3.2 Overview of Our Statistical Estimation Approach
LLM detectors are known to have unstable performance (Section 4.3). Thus, rather than trying to classify each document in the corpus and directly count the number of occurrences in this manner, we take a maximum likelihood approach. Our method has three components: training data generation, document probability distribution estimation, and computing the final estimate of the fraction of text that has been substantially modified or generated by AI. The method is summarized graphically in Figure 2. A non-graphical summary is as follows:
@@
    Collect the writing instructions given to (human) authors for the original corpus- in our case, peer review instructions. Give these instructions as prompts into an LLM to generate a corresponding corpus of AI-generated documents (Section 3.4).
@@
    Using the human and AI document corpora, estimate the reference token usage distributions $P$ and $Q$ (Section 3.5).
@@
    Verify the method’s performance on synthetic target corpora where the correct proportion of AI-generated documents is known (Section 3.6).
@@
    Based on these estimates for $P$ and $Q$, use MLE to estimate the fraction $\alpha$ of AI-generated or modified documents in the target corpus (Section 3.3).
@@ 3.4 Generating the Training Data
To generate the AI corpus, each of the reviewer instruction prompts and papers associated with the reviews in the human corpus should be fed into an AI language tool (e.g., ChatGPT), and the LLM should be prompted to generate a review. The prompts may be fed into several different LLMs to generate training data which are more robust to the choice of AI generator used. The texts output by the LLM are then collected into the AI corpus. Empirically, we found that our framework exhibits moderate robustness to the distribution shift of LLM prompts. As discussed in Appendix D.3, training with one prompt and testing with a different prompt still yield accurate validation results (see Supp. Figure 15).
@@ 3.5 Estimating $P$ and $Q$ from Data
and similarly for $Q$. Recall that our token vocabulary $V$ (defined in Section 3.1) consists of all adjectives, so the product over $t\not\in x$ means the product only over all adjectives $t$ which were not in the document or sentence $x$.
@@ 3.6 Validating the Method
We partition the human and AI corpora into two disjoint parts, with 80% used for training and 20% used for validation. We use the training partitions of the human and AI corpora to estimate $P$ and $Q$ as described above. To validate the system’s performance, we do the following:
@@ 4.1 Data
We collect review data for all major ML conferences available on OpenReview, including ICLR, NeurIPS, CoRL, and EMNLP, as detailed in Table 1. Additional information on the datasets can be found in Appendix C.
@@ 5 Discussion
While our study primarily utilized GPT-4 for generating AI texts, as GPT-4 has been one of the most widely used LLMs for long-context content generation, we found that our results are robust on the use of alternative LLMs such as GPT-3.5. For example, the model trained with only GPT-3.5 data provides consistent estimation results and findings, and demonstrates the ability to generalize, accurately detecting GPT-4 as well (see Supp. Table 28 and 29). We recommend that future practitioners select the LLM that most closely mirrors the language model likely used to generate their target corpus, reflecting actual usage patterns at the time of creation. In addition, the approximations made to the review generating process in Section 3 in order to make estimation of the review likelihood tractable introduce an additional source of error, as does the temporal distribution shift in token frequencies due to, e.g., changes in topics, reviewers, etc.
@@
We emphasize here that we do not wish to pass a value judgement or claim that the use of AI tools for review papers is necessarily bad or good. We also do not claim (nor do we believe) that many reviewers are using ChatGPT to write entire reviews outright. Our method does not constitute direct evidence that reviewers are using ChatGPT to write reviews from scratch. For example, it is possible that a reviewer may sketch out several bullet points related to the paper and uses ChatGPT to formulate these bullet points into paragraphs. In this case, it is possible for the estimated $\alpha$ to be high; indeed our results in Section $\S$ 4.6 is consistent with this mode of using LLM to substantially modify and flesh out reviews. For transparency and accountability, it is important to estimate how much of the final text might be generated by AI. We hope our data and analyses can provide fodder for constructive discussions by the community.
@@ Acknowledgments
We thank Dan Jurafsky, Diyi Yang, Yongchan Kwon, Federico Bianchi, and Mert Yuksekgonul for their helpful comments and discussions. J.Z. is supported by the National Science Foundation (CCF 1763191 and CAREER 1942926), the US National Institutes of Health (P30AG059307 and U01MH098953) and grants from the Silicon Valley Foundation and the Chan-Zuckerberg Initiative. D.F. and H.L. are supported by the National Science Foundation (2244804 and 2022435) and the Stanford Institute for Human-Centered Artificial Intelligence (HAI).
@@ Appendix B Top 100 adverbs that are disproportionately used more frequently by AI
Appendix C Dataset Details
@@ D.1 Validation Accuracy Tables
We also trained a separate model for Nature family journals using official review data for papers accepted between 2021-09-13 and 2022-08-03. We validated the model’s accuracy on reviews for papers accepted between 2022-08-04 and 2022-11-29 (Figure 3, Nature Portfolio ’22, Supp. Table 6).
@@
Table 5: Performance validation of our model across ICLR ’23, NeurIPS ’22, and CoRL ’22 reviews (all predating ChatGPT’s launch), using a blend of official human and LLM-generated reviews. Our algorithm demonstrates high accuracy with less than 2.4% prediction error in identifying the proportion of LLM reviews within the validation set. This table presents the results data for Figure. 3.
@@
Table 6: Performance validation of our model across Nature family journals (all predating ChatGPT’s launch), using a blend of official human and LLM-generated reviews. This table presents the results data for Figure. 3.
@@ D.2 Main Results Tables
Table 7: Temporal trends of ML conferences in the $\alpha$ estimate on official reviews using adjectives. $\alpha$ estimates pre-ChatGPT are close to 0, and there is a sharp increase after the release of ChatGPT. This table presents the results data for Figure. 4.
@@
Table 8: Temporal trends of the Nature family journals in the $\alpha$ estimate on official reviews using adjectives. Contrary to the ML conferences, the Nature family journals did not exhibit a significant increase in the estimated $\alpha$ values following ChatGPT’s release, with pre- and post-release $\alpha$ estimates remaining within the margin of error for the $\alpha=0$ validation experiment. This table presents the results data for Figure. 4.
@@ Appendix E LLM prompts used in the study
Appendix F Theoretical Analysis on the Sample Size
Theorem F.1.
@@ Appendix F Theoretical Analysis on the Sample Size
By Lemma F.2, we have $|\mathcal{L}(\hat{\alpha})-\mathcal{L}(\alpha^{*})|\leq O(\sqrt{\log(1/\delta)/n})$ with probability $1-\delta$. Thus, we have
@@
Now note that $\mathcal{L}(\alpha)=EZ_{i}$ and $\hat{\mathcal{L}}(\alpha)=\frac{1}{n}\sum_{i=1}^{n}Z_{i}$. We can apply Lemma F.3 to obtain that with probability $1-\epsilon$,

### 15 Jun 2024 v2

@@ 2 Related Work
Text watermarking introduces a method to detect AI-generated text by embedding unique, algorithmically-detectable signals -known as watermarks- directly into the text. Early watermarking approaches modify pre-existing text by leveraging synonym substitution (Chiang et al., 2003; Topkara et al., 2006b), syntactic structure restructuring (Atallah et al., 2001; Topkara et al., 2006a), or paraphrasing (Atallah et al., 2002). Increasingly, scholars have focused on integrating a watermark directly into an LLM’s decoding process. Kirchenbauer et al. (2023) split the vocabulary into red-green lists based on hash values of previous n-grams and then increase the logits of green tokens to embed the watermark. Zhao et al. (2023) use a global red-green list to enhance robustness. Hu et al. (2023b); Kuditipudi et al. (2023); Wu et al. (2023) study watermarks that preserve the original token probability distributions. Meanwhile, semantic watermarks (Hou et al., 2023; Fu et al., 2023; Liu et al., 2023) using input sequences to find semantically related tokens and multi-bit watermarks (Yoo et al., 2023; Fernandez et al., 2023) to embed more complex information have been proposed to improve certain conditional generation tasks. However, watermarking requires the involvement of the model or service owner, such as OpenAI, to implant the watermark. Concerns have also been raised regarding the potential for watermarking to degrade text generation quality and to compromise the coherence and depth of LLM responses (Singh & Zou, 2023). In contrast, our framework operates independently of the model or service owner’s intervention, allowing for the monitoring of AI-modified content without requiring their adoption.
@@ 3.2 Overview of Our Statistical Estimation Approach
LLM detectors are known to have unstable performance (Section $\S$ 4.3). Thus, rather than trying to classify each document in the corpus and directly count the number of occurrences in this manner, we take a maximum likelihood approach. Our method has three components: training data generation, document probability distribution estimation, and computing the final estimate of the fraction of text that has been substantially modified or generated by AI. The method is summarized graphically in Figure 2. A non-graphical summary is as follows:
@@
    Collect the writing instructions given to (human) authors for the original corpus- in our case, peer review instructions. Give these instructions as prompts into an LLM to generate a corresponding corpus of AI-generated documents (Section $\S$ 3.4).
@@
    Using the human and AI document corpora, estimate the reference token usage distributions $P$ and $Q$ (Section $\S$ 3.5).
@@
    Verify the method’s performance on synthetic target corpora where the correct proportion of AI-generated documents is known (Section $\S$ 3.6).
@@
    Based on these estimates for $P$ and $Q$, use MLE to estimate the fraction $\alpha$ of AI-generated or modified documents in the target corpus (Section $\S$ 3.3).
@@ 3.4 Generating the Training Data
To generate the AI corpus, we prompt the LLM to generate a review given a paper. The texts output by the LLM are then collected into the AI corpus. Empirically, we found that our framework exhibits moderate robustness to the distribution shift of LLM prompts. As discussed in Appendix D.3, training with one prompt and testing with a different prompt still yield accurate validation results (see Supp. Figure 15).
@@ 3.5 Estimating $P$ and $Q$ from Data
and similarly for $Q$. Recall that our token vocabulary $V$ (defined in Section $\S$ 3.1) consists of all adjectives, so the product over $t\not\in x$ means the product only over all adjectives $t$ which were not in the document or sentence $x$.
We validated both approaches using either a document or a sentence as the unit of $x$, and both performed well (Appendix D.8). We used a sentence as our main unit for estimates, as sentences perform slightly better.
@@ 3.6 Validating the Method
We use the training partitions of the human and AI corpora to estimate $P$ and $Q$ as described above. To validate the system’s performance, we do the following:
@@ 4.1 Data
We collect review data for all major ML conferences available on OpenReview, including ICLR, NeurIPS, CoRL, and EMNLP, as detailed in Table 1. The Nature portfolio dataset encompasses 15 journals within the Nature portfolio, such as Nature, Nature Biomedical Engineering, Nature Human Behaviour, and Nature Communications. Additional information on the datasets can be found in Appendix G.
@@ 5 Discussion
While our study focused on ChatGPT, which dominates the generative AI market with 76% of global internet traffic in the category (Van Rossum, 2024), we acknowledge that there are other diverse LLMs used for generating or rephrasing text. However, recent studies have found that ChatGPT substantially outperforms other LLMs, including Bard, in the reviewing of scientific papers or proposals (Liang et al., 2023c; Liu & Shah, 2023). We also found that our results are robust on the use of alternative LLMs such as GPT-3.5. For example, the model trained with only GPT-3.5 data provides consistent estimation results and findings, and demonstrates the ability to generalize, accurately detecting GPT-4 as well (see Supp. Table 28 and 29). However, we acknowledge that our framework’s effectiveness may vary depending on the specific LLM used, and future practitioners should select the LLM that most closely mirrors the language model likely used to generate their target corpus, reflecting actual usage patterns at the time of creation.
Our findings are primarily based on datasets from major ML conferences (ICLR, NeurIPS, CoRL, EMNLP) and Nature Family Journals spanning 15 distinct journals across different disciplines such as medicine, biology, chemistry, and environmental sciences. While this demonstrates the applicability of our framework beyond these domains, further experimentation may be required to fully establish its generalizability to an even wider range of fields and publication venues. Factors such as field-specific writing styles and the prevalence of AI use could influence the effectiveness of our approach.
Moreover, the prompting techniques used in our study to simulate the process of revising, expanding, paraphrasing, and proofreading review texts (Section $\S$ 4.5) have limitations. The prompts we employed were designed based on our understanding of common practices, but they may not capture the full range of techniques used by reviewers or AI assistants. We emphasize that these techniques should be interpreted as a best-effort approximation rather than a definitive representation of how AI is used for review text modifications.
Although our validation experiments used real reviews from prior years, which included a significant fraction of non-native speaker-written texts, and our results remained accurate, we recognize that substantial shifts in the non-native speaker population over time could still impact the accuracy of our estimates (Liang et al., 2023a). Future research should investigate the impact of evolving non-native speaker populations on the robustness of our framework.
In addition, the approximations made to the review generating process in Section $\S$ 3 in order to make estimation of the review likelihood tractable introduce an additional source of error, as does the temporal distribution shift in token frequencies due to, e.g., changes in topics, reviewers, etc.
@@
We emphasize here that we do not wish to pass a value judgement or claim that the use of AI tools for review papers is necessarily bad or good. We also do not claim (nor do we believe) that many reviewers are using ChatGPT to write entire reviews outright. Our method does not constitute direct evidence that reviewers are using ChatGPT to write reviews from scratch. For example, it is possible that a reviewer may sketch out several bullet points related to the paper and uses ChatGPT to formulate these bullet points into paragraphs. In this case, it is possible for the estimated $\alpha$ to be high; indeed our results in Appendix 4.6 is consistent with this mode of using LLM to substantially modify and flesh out reviews.
To enhance transparency and accountability, future work should focus on applying and extending our framework to estimate the extent of AI-generated text across various domains, including but not limited to peer review. We believe that our data and analyses can serve as a foundation for constructive discussions and further research by the community, ultimately contributing to the development of robust guidelines and best practices for the ethical use of generative AI.
Code Availability
@@ Acknowledgments
We thank Christopher D. Manning, Dan Jurafsky, Diyi Yang, Yongchan Kwon, Federico Bianchi, and Mert Yuksekgonul for their helpful comments and discussions. J.Z. is supported by the National Science Foundation (CAREER 1942926) and grants from the Chan-Zuckerberg Initiative and Stanford HAI. D.M. and H.L. are supported by the National Science Foundation (2244804 and 2022435) and the Stanford Institute for Human-Centered Artificial Intelligence (HAI).
@@ Appendix B Top 100 adverbs that are disproportionately used more frequently by AI
Appendix C Additional Details on Major ML Conferences Reviewer Confidence Scale
@@ D.1 Validation Accuracy Tables
We also trained a separate model for Nature family journals using official review data for papers accepted between 2021-09-13 and 2022-08-03. We validated the model’s accuracy on reviews for papers accepted between 2022-08-04 and 2022-11-29 (Figure 3, Nature Portfolio ’22, Table 6).
@@
Table 5: Performance validation of our model across ICLR ’23, NeurIPS ’22, and CoRL ’22 reviews (all predating ChatGPT’s launch), using a blend of official human and LLM-generated reviews. Our algorithm demonstrates high accuracy with less than 2.4% prediction error in identifying the proportion of LLM reviews within the validation set. This table presents the results data for Figure  3.
@@
Table 6: Performance validation of our model across Nature family journals (all predating ChatGPT’s launch), using a blend of official human and LLM-generated reviews. This table presents the results data for Figure  3.
@@ D.2 Main Results Tables
Table 7: Temporal trends of ML conferences in the $\alpha$ estimate on official reviews using adjectives. $\alpha$ estimates pre-ChatGPT are close to 0, and there is a sharp increase after the release of ChatGPT. This table presents the results data for Figure  4.
@@
Table 8: Temporal trends of the Nature family journals in the $\alpha$ estimate on official reviews using adjectives. Contrary to the ML conferences, the Nature family journals did not exhibit a significant increase in the estimated $\alpha$ values following ChatGPT’s release, with pre- and post-release $\alpha$ estimates remaining within the margin of error for the $\alpha=0$ validation experiment. This table presents the results data for Figure  4.
@@ D.12 Factors that Correlate With Estimated LLM Usage
Here we chose to focus on ChatGPT because it is by far the most popular in general usage. According to a comprehensive analysis by FlexOS in early 2024, ChatGPT dominates the generative AI market, with 76% of global internet traffic in the category. Bing AI follows with 16%, Bard with 7%, and Claude with 1% (Van Rossum, 2024). Recent studies have also found that GPT-4 substantially outperforms other LLMs, including Bard, in the reviewing of scientific papers or proposals (Liu & Shah, 2023).
@@ Appendix E LLM prompts used in the study
Your task is to proofread the provided sentence for grammatical accuracy. Ensure that the corrections introduce minimal distortion to the original content.
Figure 31: Example prompt for proofreading.
Appendix F Additional Information on LLM Parameter Settings
We used the snapshot of GPT-4 from June 13th, 2023 (gpt-4-0613), for our experiments because this is the exact version of ChatGPT that was available during the peer review process of ICLR 2024, NeurIPS 2023, EMNLP 2023, and CoRL 2023.
Regarding the parameter settings, during our experiments, we set the decoding temperature to 1.0 and the maximum decoding length to 2048. We set the Top P hyperparameter to 1.0 and both frequency penalty and presence penalty to 0.0. Additionally, we did not configure any stop sequences during decoding.
Appendix G Additional Dataset Information
All the data are publicly available. For the machine learning conferences, we accessed peer review data through the official OpenReview API (https://docs.openreview.net/reference/api-v2), specifically the /notes endpoint. Each review contains an average of 25.94 sentences. For the Nature portfolio dataset, we developed a custom web scraper using python to access the article pages of 15 journals from the Nature portfolio, extracting peer reviews from papers accepted between 2019 and 2023. Each review in the Nature dataset comprises an average of 37.03 sentences.
The Nature portfolio dataset encompasses the following 15 Nature journals: Nature, Nature Communications, Nature Ecology & Evolution, Nature Structural & Molecular Biology, Nature Cell Biology, Nature Human Behaviour, Nature Immunology, Nature Microbiology, Nature Biomedical Engineering, Communications Earth & Environment, Communications Biology, Communications Physics, Communications Chemistry, Communications Materials, and Communications Medicine. To create this dataset, we systematically accessed the web pages of the selected Nature portfolio journals, extracting peer reviews from papers accepted between 2019 and 2023. In total, our dataset comprises 25,382 peer reviews from 10,242 papers. We chose to focus on the Nature family journals for our baseline dataset due to their reputation for publishing high-quality, impactful research across multiple disciplines.
Our framework breaks reviews down into a list of sentences, and the parameterization operates at the sentence level. We consider all sentences with 2 or more words and did not set a maximum limit for the number of words in a sentence. If reviewers leave a section blank, no sentences from that section are added to the corpus.
Table 31: Human Peer Reviews Data from Nature Family Journals.
Journal Post ChatGPT Data Split # of Papers # of Official Reviews
Nature Portfolio 2022 (random split subset) Before Training 1,189 3,341
Nature Portfolio 2019 Before Validation 2,141 4,394
Nature Portfolio 2020 Before Validation 2,083 4,736
Nature Portfolio 2021 Before Validation 2,129 5,264
Nature Portfolio 2022 Before Validation 511 1,447
Nature Portfolio 2022-2023 After Inference 2,189 6,200
Ethics Considerations About LLM Analysis for Public Conferences
The use of peer review data for research purposes raises important ethical considerations around reviewer consent, data licensing, and responsible use (Dycke et al., 2023). While early datasets have enabled valuable research, going forward, it is critical that the community establishes clear best practices for the ethical collection and use of peer review data.
OpenReview is a science communication initiative which aims to make the scientific process more transparent. Authors and peer reviewers agree to make their reviews public upon submission. In our work, we accessed this publicly available, anonymous peer review data through the public OpenReview API and confirm that we have complied with their terms of use.
Efforts such as the data donation initiative at ACL Rolling Review (ARR), which requires explicit consent from authors and reviewers and provides clear data licenses (Dycke et al., 2022), provide a promising model for the future. A key strength of our proposed framework is that it operates at the population level and only outputs aggregate statistics, without the need to perform inference on individual reviews. This helps protect the anonymity of reviewers and mitigates the risk of de-anonymization based on writing style, which is an important consideration when working with peer review data.
We have aimed to use the available data responsibly and ethically in our work. We also recognize the importance of developing robust community norms around the appropriate collection, licensing, sharing, and use of peer review datasets.
Appendix H Additional Results on LLaMA-2 Chat (70B), and Claude 2.1
We have added additional validation experiments to test two additional models other than GPT-4: LLaMA-2 Chat (70B) and Claude-2.1. We used the same training and validation setup as our paper. We trained the estimator on ICLR 2018-2022 data, and performed the validation of different alphas on ICLR 2023 data. In the first experiment, we trained an estimator using data generated by LLaMA-2 Chat (70B). In the second experiment, we trained an estimator using data generated by Claude 2.1. As shown in the two results tables, our framework predicts the proportion of AI data (i.e., alpha) very well.
Table 32: Validation Data Source Performance Comparison for LLaMA-2 Chat (70B).
No. Validation Data Source Ground Truth $\alpha$ Estimated
$\alpha$ CICI ($\pm$)
(1) ICLR 2023 (LLaMA-2 Chat (70B)) 0% 2.8% 0.5%
(2) ICLR 2023 (LLaMA-2 Chat (70B)) 2.5% 5.3% 0.5%
(3) ICLR 2023 (LLaMA-2 Chat (70B)) 5% 7.6% 0.5%
(4) ICLR 2023 (LLaMA-2 Chat (70B)) 7.5% 9.9% 0.5%
(5) ICLR 2023 (LLaMA-2 Chat (70B)) 10% 12.2% 0.6%
(6) ICLR 2023 (LLaMA-2 Chat (70B)) 12.5% 14.6% 0.6%
(7) ICLR 2023 (LLaMA-2 Chat (70B)) 15% 17% 0.6%
(8) ICLR 2023 (LLaMA-2 Chat (70B)) 17.5% 19.2% 0.6%
(9) ICLR 2023 (LLaMA-2 Chat (70B)) 20% 21.6% 0.7%
(10) ICLR 2023 (LLaMA-2 Chat (70B)) 22.5% 24% 0.7%
(11) ICLR 2023 (LLaMA-2 Chat (70B)) 25% 26.3% 0.7%
Table 33: Validation Data Source Performance Comparison for Claude-2.1.
No. Validation Data Source Ground Truth $\alpha$ Estimated
$\alpha$ CICI ($\pm$)
(1) ICLR 2023 (Claude-2.1) 0% 4.2% 0.7%
(2) ICLR 2023 (Claude-2.1) 2.5% 6.5% 0.8%
(3) ICLR 2023 (Claude-2.1) 5% 8.7% 0.8%
(4) ICLR 2023 (Claude-2.1) 7.5% 10.9% 0.8%
(5) ICLR 2023 (Claude-2.1) 10% 13.2% 0.9%
(6) ICLR 2023 (Claude-2.1) 12.5% 15.4% 0.9%
(7) ICLR 2023 (Claude-2.1) 15% 17.6% 0.9%
(8) ICLR 2023 (Claude-2.1) 17.5% 19.9% 0.9%
(9) ICLR 2023 (Claude-2.1) 20% 22.1% 0.9%
(10) ICLR 2023 (Claude-2.1) 22.5% 24.4% 0.9%
(11) ICLR 2023 (Claude-2.1) 25% 26.5% 0.9%
Appendix I Theoretical Analysis on the Sample Size
Theorem I.1.
@@ Appendix I Theoretical Analysis on the Sample Size
By Lemma I.2, we have $|\mathcal{L}(\hat{\alpha})-\mathcal{L}(\alpha^{*})|\leq O(\sqrt{\log(1/\delta)/n})$ with probability $1-\delta$. Thus, we have
@@
Now note that $\mathcal{L}(\alpha)=EZ_{i}$ and $\hat{\mathcal{L}}(\alpha)=\frac{1}{n}\sum_{i=1}^{n}Z_{i}$. We can apply Lemma I.3 to obtain that with probability $1-\epsilon$,